# algorithmic

### v3

@@ Abstract
Keywords Kalman filtering  $\cdot$ non-smooth regularization  $\cdot$ online learning  $\cdot$ parameter estimation  $\cdot$ adaptive control  $\cdot$ neural networks.
@@ 1 Introduction
Online learning of nonlinear parametric models is of paramount importance in several domains, including model-based adaptive control and real-time estimation of unmeasured variables. Typically, parametric models derived from physics [1] or black-box [2] structures are identified offline on training data, then directly deployed and used without any further updates. On the other hand, further adapting the model online can significantly improve its predictive capabilities [3], especially when the phenomenon we are modeling changes over time, and allows for smaller model structures that adapt to varying operating conditions, unlike single, overall models trained offline to cover all conditions.
@@
A vast literature exists for online learning [4], and several approaches using first or second order gradient information are suitable for real-time model adaptation. Examples of algorithms exploiting first-order information are online gradient descent (OGD) and its regularized alternatives, i.e., follow the regularized leader (FTRL) and online mirror descent (OMD) [5], which are usually fast at execution but slow at convergence. Second-order gradient information can be used to improve the convergence speed, such as in the online Newton step (ONS), or in the well known extended Kalman filter (EKF) [6] which has been proved to be very effective in online model adaptation by treating the parameters as constant states to be estimated [7, 8, 9].
@@
While FTRL and OMD, as well as modified versions of the standard EKF [8], can effectively deal with smooth regularization terms, non-smooth regularizers are often required when learning models. Non-smooth $\ell_{0}$ or $\ell_{1}$ penalties are used to induce sparsity in the model [10], and group-Lasso terms to remove entire parts of the model, such as neurons of a neural network [11]. Indeed, obtaining compact models is particularly important for embedded applications, such as model predictive control, where the real-time numerical complexity of the controller depends directly on the complexity of the prediction model. In addition, indicator functions of feasible sets are another example of non-smooth penalties that are used to impose constraints on model coefficients, such as known bounds on certain unknown physical parameters [12].
@@
Several approaches have been proposed for online learning under non-smooth regularization. Online ADMM (alternating direction method of multipliers) [13] allows handling quite general non-smooth regularizers, while EKF can be modified to deal with $\ell_{1}$-regularization by treating it as a special limit case of a smooth regularizer [8]. While the main limitation of online ADMM is its convergence speed, the main limitation of EKF is that it is not directly suitable for dealing with general non-smooth regularization terms, such as $\ell_{0}$ regularization, group-Lasso penalties, and indicator functions of feasible sets. The main contribution of this work is to show interesting connections and similarities between these two approaches, developing an extension of EKF that inherits the advantages from both, i.e., the convergence speed of EKF and the capability to handle general non-smooth regularization terms of online-ADMM.
@@
The proposed method consists of a simple and computationally efficient modification of the EKF algorithm by intertwining updates based on online measurements and output prediction errors with updates related to ADMM iterations. This modification allows EKF to deal with a broad class of non-smooth regularization terms for which ADMM is applicable, including $\ell_{0}$/$\ell_{1}$ penalties, group-Lasso and indicator functions of simple sets. For linear time-varying models and convex regularization terms, we provide a sublinear regret bound that proves the proper behavior of the resulting online learning strategy. The proposed method is computationally efficient and numerically robust, making it especially appealing for embedded adaptive control applications.
@@ 1.1 Notation
Given a matrix $A\in\mathbb{R}^{m\times n}$, $A_{i:}$ denotes the ith row of A, $A_{:j}$ its $j$th column, $A_{ij}$ its $(i,j)$th entry. Given a vector $v\in\mathbb{R}^{n}$ we denote by $v^{m}$ a measurement of $v$, by $\lVert v\rVert_{1}=\sum_{i=1}^{n}|v_{i}|$ the 1-norm of $v$ and by $\lVert v\rVert_{0}$ its 0-norm, which is defined as the number of non-zero elements in the vector. Given a symmetric positive semidefinite matrix $P=P^{T}\succeq 0$, $P\in\mathbb{R}^{n\times n}$, we denote by $\lVert v\rVert^{2}_{P}$ the quadratic form $v^{T}Pv$. Further, we denote by $\hat{v}_{i|j}$ the estimate of vector $v$ at instant $i$ given all information up to instant $j$, and by $P_{i|j}$ the corresponding covariance matrix. $v\sim\mathcal{N}(\mu,P)$ and $v\sim\mathcal{U}_{[v_{min},v_{max}]}$ denote that v was randomly generated from a normal distribution with mean $\mu$ and covariance $P$ or from a uniform distribution in the interval $[v_{min},v_{max}]$, respectively.
@@ 2 EKF for online model learning
with $y_{k}=y_{k}^{nl}-h_{k}(\overline{x}_{k})+C_{k}\overline{x}_{k}$. The classical Kalman filter [6] can be used to estimate the state in (3), i.e., to learn the parameters $x_{k}$ recursively. Given $\hat{x}_{0|-1},P_{0|-1}$ we perform the following iterations for $k=0,\dots,N-1$:
@@
As shown in [14], the state estimates $\hat{x}_{k|k},\hat{x}_{k+1|k}$ generated by the Kalman filter (4) are part of the optimizer of the following optimization problem
$\hat{x}_{0|k},\dots,\hat{x}_{k|k},\hat{x}_{k+1|k}=\operatorname*{arg\,min}_{x_{0},\dots,x_{k},x_{k+1}}\left\lVert x_{0}-\hat{x}_{0|-1}\right\rVert^{2}_{P_{0|-1}^{-1}}+\sum_{i=0}^{k}\left\lVert y_{i}^{m}-C_{i}x_{i}\right\rVert^{2}_{R_{i}^{-1}}+\left\lVert x_{i+1}-x_{i}\right\rVert^{2}_{Q_{i}^{-1}}.$ (5)
@@
$\hat{x}_{k|k}=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}~{}$ (6a)
$\hat{x}_{k|k},\hat{x}_{k+1|k}=\operatorname*{arg\,min}_{x_{k},x_{k+1}}\left\lVert x_{k}-\hat{x}_{k|k}\right\rVert^{2}_{P_{k|k}^{-1}}+\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{Q_{k}^{-1}}~{}$ (6b)
@@ 3 EKF under non-smooth regularization
In order to regularize the model, we modify the classical iterations (4) by changing the minimization in (6a) to
$\hat{x}_{k|k}=\operatorname*{arg\,min}_{x_{k}}\frac{1}{2}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+g(x_{k})$ (7)
@@
$\hat{x}_{k|k},\nu^{\star}=\operatorname*{arg\,min}_{(x_{k},\nu)\in\mathcal{S}}\frac{1}{2}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+g(\nu)$ (8)
which can be solved by executing the following scaled ADMM iterations [12]:
$\hat{x}_{k|k}^{t+1}$ $=$ $\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+\rho\left\lVert x_{k}-\nu^{t}+w^{t}\right\rVert_{2}^{2}$ (9a)
@@
for $t=0,\dots,n_{a}-1$, where $\rho>0$ is a hyper-parameter to be calibrated and “prox” is the proximal operator [15]. As shown in [12], in the convex case, the ADMM iterations (9a)–(9c) converge to the optimizer of (8) as $n_{a}\rightarrow\infty$, and often converge to a solution of acceptable accuracy within a few tens of iterations. Iteration (9c) is straightforward to compute; iteration (9b) can be solved explicitly and efficiently with complexity $\mathcal{O}(n_{x})$ for a wide range of non-smooth and non-convex regularization functions $g$, such as $g(x)=\|x\|_{0}$, $g(x)=\|x\|_{1}$, and the indicator function $g(x)=0$ if $x_{\rm min}\leq x\leq x_{\rm max}$ or +$\infty$+ ∞ otherwise [15]. Iteration (9a) can be rewritten as
@@
Algorithm 1 summarizes the proposed extension of EKF with ADMM iterations (EKF-ADMM). The algorithm returns the estimate $\hat{x}_{k|k}$ of the parameter vector $x$ obtained after processing $N$ measurements. It also returns the last value of $\nu$, which could be used as an alternative estimate of $x$ too; for example, in case $g$ is the indicator function of a constraint set, $\nu$ would be guaranteed to be feasible. Note that the dual vector $w$ is not reset at each EKF iteration $k$; it is used as a warm start for the next $n_{a}$ ADMM iterations at step $k+1$, as the solutions $\hat{x}_{k|k}$ at consecutive time instants $k$ are usually similar.
@@ 3.1 Computational complexity
$\hat{x}_{k|k}^{t+1}=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+\rho\sum_{i=1}^{n_{x}}\left\lVert x_{k,i}-\nu^{t}_{i}+w^{t}_{i}\right\rVert_{2}^{2}$ (11)
which highlights the separation of the contributions of the true measurements $y_{k}^{m}$ and of the fake regularization measurements $\nu^{t}-w^{t}$ in the correction step; moreover, we can process the measurements $\nu^{t}_{i}-w^{t}_{i}$ separately one by one. This allows designing a computationally more efficient and numerically robust version of the proposed EKF-ADMM algorithm, as the correction due to the true measurements $y_{k}^{m}$ can be performed only once, instead of $n_{a}$ times, as it does not change with $t$. Moreover, there is no need for any matrix inversion when processing the fake measurements, as by processing them one by one, the matrix inversion required to compute the Kalman gain becomes a simple division, since each measurement is just a scalar value. Assuming a complexity $\mathcal{O}(n_{x})$ for evaluating the proximal operator, EKF-ADMM has complexity $\mathcal{O}(n_{x}^{3}+n_{a}n_{x}^{2})$, which is the same order of the full EKF for general state estimation. Moreover, EKF-ADMM has the same number of Jacobian matrices evaluations than the classical EKF, which is usually the most time-consuming part in case $x$ represents the weights and bias terms of a neural network model to learn. Summarizing, the proposed approach is computationally efficient and, if the Kalman filter is implemented using numerically robust factored or square-root modifications [16], the method is appealing for embedded applications.
@@ 4 Regret analysis
where $C_{k}$ are now given time-varying matrices for $k=0,1,\ldots,N-1$, and convex regularization terms $g$. In particular, we want to evaluate the ability of the algorithm to solve the optimization problem $\min_{x}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$ online, where $f_{k}(x)=\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x\right\rVert^{2}_{R_{k}^{-1}}$, via the following two regret functions $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(\nu_{k}))-\min_{x,\nu\in\mathcal{S}}\sum_{k=0}^{N-1}(f_{k}(x)+g(\nu))$ and $R_{c}(N)=\sum_{k=0}^{N-1}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}$, where, to simplify the notation, we have defined $x_{k}=\hat{x}_{k|k-1},P_{k}=P_{k|k-1},\forall k=0,1,\ldots,N-1$. Notice that $R_{f}(N)$ quantifies the loss we suffer by learning the model online instead of solving it in a batch way given all $N$ measurements, while $R_{c}(N)$ quantifies the violation of the constraint $x=\nu$. To ensure a proper behavior of EKF-ADMM, we want to prove a sublinear regret bound for both, i.e., $R_{f}(N)\leq\mathcal{O}(\sqrt{N})$ and $R_{c}(N)\leq\mathcal{O}(\sqrt{N})$ [13].
EKF-ADMM is a generalization of the online ADMM method proposed in [13], in which a sublinear regret bound is derived for the case $n_{a}=1$ and $P_{k}^{-1}=P^{-1}\succ 0,\forall k$, while, more recently, in [17] a sublinear regret bound has been derived for the case $n_{a}=1$ and $P_{k}^{-1}\succeq P_{k+1}^{-1},\forall k$. Here we will provide a sublinear regret in the case $n_{a}=1$ and $P_{k}^{-1}=P_{k+1}^{-1},\forall k\geq k_{n}\ll N$, which is a reasonable assumption as the EKF covariance matrix, when estimating the parameters of a model, usually has a transient and then reaches a steady-state value. By assuming $n_{a}=1$ and $P_{k}^{-1}=P_{k+1}^{-1},\forall k\geq k_{n}\ll N$, Algorithm 1 can be equivalently rewritten as in Algorithm 2.
@@
[…]
     $\nu_{k+1}\leftarrow&\operatorname*{arg\,min}_{\nu}g(\nu)+w_{k}^{T}(x_{k+1}-\nu)+\\ &+\frac{\rho}{2}\left\lVert x_{k+1}-\nu\right\rVert^{2}=\text{prox}_{\frac{g}{\rho}}(x_{k+1}+\frac{w_{k}}{\rho})\end{aligned}$
@@
The following Theorem 4.1 is an extension of [13, Theorem 4], and provides conditions for sublinear regret bounds of Algorithm 2 in the case of a linear time-varying model (12) and convex regularization function $g$.
Theorem 4.1
@@
In general, as proved in [17], Theorem 4.1 holds with $M_{k_{n}}=0$ whenever $P_{k}^{-1}\succeq P_{k+1}^{-1},\forall k$. Intuitively, this means that for online model adaptation we need to limit the importance of the previous samples to promptly adapt the model to changes and therefore bound the regret function. This can be accomplished, for example, using the EKF with a proper forgetting factor [9].

### v4

# Online Learning of Nonlinear Parametric Models under Non-smooth Regularization using EKF and ADMM

@@ Abstract
keywords:
Kalman filtering; non-smooth regularization; online learning; parameter estimation; adaptive control; neural networks.
thanks: The research work of Lapo Frascati has been financially supported by ODYS S.r.l. The research work of Alberto Bemporad has been funded by the European Union (ERC Advanced Research Grant COMPACT, No. 101141351). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ 1 Introduction
Online learning of nonlinear parametric models is of paramount importance in several domains, including model-based adaptive control and real-time estimation of unmeasured variables. Typically, parametric models derived from physics [bib9] or black-box [bib8] structures are identified offline on training data, then directly deployed and used without any further updates. On the other hand, further adapting the model online can significantly improve its predictive capabilities [bib10], especially when the phenomenon we are modeling changes over time, and allows for smaller model structures that adapt to varying operating conditions, unlike single, overall models trained offline to cover all conditions.
@@
A vast literature exists for online learning [bib7], and several approaches using first or second order gradient information are suitable for real-time model adaptation. Examples of algorithms exploiting first-order information are online gradient descent (OGD) and its regularized alternatives, i.e., follow the regularized leader (FTRL) and online mirror descent (OMD) [bib20], which are usually fast at execution but slow at convergence. Second-order gradient information can be used to improve the convergence speed, such as in the online Newton step (ONS), or in the well known extended Kalman filter (EKF) [bib14] which has been proved to be very effective in online model adaptation by treating the parameters as constant states to be estimated [bib1, bib3].
@@
While FTRL and OMD, as well as modified versions of the standard EKF [bib1], can effectively deal with smooth regularization terms, non-smooth regularizers are often required when learning models. Non-smooth $\ell_{0}$ or $\ell_{1}$ penalties are used to induce sparsity in the model [bib25], and group-Lasso terms to remove entire parts of the model, such as neurons of a neural network [bib26]. Indeed, obtaining compact models is particularly important for embedded applications, such as model predictive control, where the real-time numerical complexity of the controller depends directly on the complexity of the prediction model. In addition, indicator functions of feasible sets are another example of non-smooth penalties that are used to impose constraints on model coefficients, such as known bounds on certain unknown physical parameters [bib5].
@@
Several approaches have been proposed for online learning under non-smooth regularization. Online ADMM (alternating direction method of multipliers) [bib2] allows handling quite general non-smooth regularizers, while EKF can be modified to deal with $\ell_{1}$-regularization by treating it as a special limit case of a smooth regularizer [bib1]. While the main limitation of online ADMM is its convergence speed, the main limitation of EKF is that it is not directly suitable for dealing with general non-smooth regularization terms, such as $\ell_{0}$ regularization, group-Lasso penalties, and indicator functions of feasible sets. The main contribution of this work is to show interesting connections and similarities between these two approaches, developing an extension of EKF that inherits the advantages from both, i.e., the convergence speed of EKF and the capability to handle general non-smooth regularization terms of online-ADMM.
@@
The proposed method consists of a simple and computationally efficient modification of the EKF algorithm by intertwining updates based on online measurements and output prediction errors with updates related to ADMM iterations. This modification allows EKF to deal with a broad class of non-smooth regularization terms for which ADMM is applicable, including $\ell_{0}$$\ell_{1}$ penalties, group-Lasso and indicator functions of simple sets. For linear time-varying models and convex regularization terms, we provide a sublinear regret bound that proves the proper behavior of the resulting online learning strategy. The proposed method is computationally efficient and numerically robust, making it especially appealing for embedded adaptive control applications.
@@ 1.1 Notation
Given a matrix $A\in\mathbb{R}^{m\times n}$, $A_{i:}$ denotes the ith row of A, $A_{:j}$ its $j$th column, $A_{ij}$ its $(i,j)$th entry. Given a vector $v\in\mathbb{R}^{n}$ we denote by $v^{m}$ a measurement of $v$, by $\lVert v\rVert_{1}=\sum_{i=1}^{n}|v_{i}|$ the 1-norm of $v$ and by $\lVert v\rVert_{0}$ its 0-norm, which is defined as the number of non-zero elements in the vector. Given a symmetric positive semidefinite matrix $P=P^{T}\succeq 0$, $P\in\mathbb{R}^{n\times n}$, we denote by $\lVert v\rVert^{2}_{P}$ the quadratic form $v^{T}Pv$. Further, we denote by $\hat{v}_{i|j}$ the estimate of vector $v$ at instant $i$ given all information up to instant $j$, and by $P_{i|j}$ the corresponding covariance matrix. $v\sim\mathcal{N}(\mu,P)$ and $v\sim\mathcal{U}_{[v_{min},v_{max}]}$ denote that v was randomly generated from a normal distribution with mean $\mu$ and covariance $P$ or from a uniform distribution in the interval $v_{min},v_{max}]$, respectively.
@@ 2 EKF for online model learning
with $y_{k}=y_{k}^{nl}-h_{k}(\overline{x}_{k})+C_{k}\overline{x}_{k}$. The classical Kalman filter [bib14] can be used to estimate the state in (3), i.e., to learn the parameters $x_{k}$ recursively. Given $\hat{x}_{0|-1},P_{0|-1}$ we perform the following iterations for $k=0,\dots,N-1$:
@@
As shown in [bib13], the state estimates $\hat{x}_{k|k},\hat{x}_{k+1|k}$ generated by the Kalman filter (4) are part of the optimizer of the following optimization problem
$\hat{x}_{0|k},\dots,\hat{x}_{k|k},\hat{x}_{k+1|k}=\operatorname*{arg\,min}_{x_{0},\dots,x_{k},x_{k+1}}\left\lVert x_{0}-\hat{x}_{0|-1}\right\rVert^{2}_{P_{0|-1}^{-1}}+$ (5)
$+\sum_{i=0}^{k}\left\lVert y_{i}^{m}-C_{i}x_{i}\right\rVert^{2}_{R_{i}^{-1}}+\left\lVert x_{i+1}-x_{i}\right\rVert^{2}_{Q_{i}^{-1}}.$
@@
$\hat{x}_{k|k}=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}\hskip-10.00002pt+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}$
$\hat{x}_{k|k},\hat{x}_{k+1|k}=\operatorname*{arg\,min}_{x_{k},x_{k+1}}\left\lVert x_{k}-\hat{x}_{k|k}\right\rVert^{2}_{P_{k|k}^{-1}}+$
$\hskip 120.00018pt+\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{Q_{k}^{-1}}$
@@ 3 EKF under non-smooth regularization
In order to regularize the model, we modify the classical iterations (4) by changing the minimization in (6b) to
$\hat{x}_{k|k}=$ $\operatorname*{arg\,min}_{x_{k}}\frac{1}{2}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+$ (7)
$+\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+g(x_{k})$
@@
$\hat{x}_{k|k},\nu^{\star}=$ $\operatorname*{arg\,min}_{(x_{k},\nu)\in\mathcal{S}}\frac{1}{2}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+$ (8)
$+\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+g(\nu)$
which can be solved by executing the following scaled ADMM iterations [bib5]:
$\hat{x}_{k|k}^{t+1}$ $=$ $\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+$ (9a)
$+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+\rho\left\lVert x_{k}-\nu^{t}+w^{t}\right\rVert_{2}^{2}$
@@
for $t=0,\dots,n_{a}-1$, where $\rho>0$ is a hyper-parameter to be calibrated and “prox” is the proximal operator [bib15]. As shown in [bib5], in the convex case, the ADMM iterations (9a)–(9c) converge to the optimizer of (8) as $n_{a}\rightarrow\infty$, and often converge to a solution of acceptable accuracy within a few tens of iterations. Iteration (9c) is straightforward to compute; iteration (9b) can be solved explicitly and efficiently with complexity $\mathcal{O}(n_{x})$ for a wide range of non-smooth and non-convex regularization functions $g$, such as $g(x)=\|x\|_{0}$, $g(x)=\|x\|_{1}$, and the indicator function $g(x)=0$ if $x_{\rm min}\leq x\leq x_{\rm max}$ or $\infty$ otherwise [bib15]. Iteration (9a) can be rewritten as
@@
Algorithm 1 summarizes the proposed extension of EKF with ADMM iterations (EKF-ADMM). The algorithm returns the estimate $\hat{x}_{k|k}$ of the parameter vector $x$ obtained after processing NN measurements. It also returns the last value of $\nu$, which could be used as an alternative estimate of $x$ too; for example, in case $g$ is the indicator function of a constraint set, $\nu$ would be guaranteed to be feasible. Note that the dual vector $w$ is not reset at each EKF iteration $k$; it is used as a warm start for the next $n_{a}$ ADMM iterations at step $k+1$, as the solutions $\hat{x}_{k|k}$ at consecutive time instants $k$ are usually similar.
@@ 3.1 Computational complexity
$\hat{x}_{k|k}^{t+1}$ $=\operatorname*{arg\,min}_{x_{k}}\left\lVert x_{k}-\hat{x}_{k|k-1}\right\rVert^{2}_{P_{k|k-1}^{-1}}+$ (11)
$+\left\lVert y_{k}^{m}-C_{k}x_{k}\right\rVert^{2}_{R_{k}^{-1}}+\rho\sum_{i=1}^{n_{x}}\left\lVert x_{k,i}-\nu^{t}_{i}+w^{t}_{i}\right\rVert_{2}^{2}$
which highlights the separation of the contributions of the true measurements $y_{k}^{m}$ and of the fake regularization measurements $\nu^{t}-w^{t}$ in the correction step; moreover, we can process the measurements $\nu^{t}_{i}-w^{t}_{i}$ separately one by one. This allows designing a computationally more efficient and numerically robust version of the proposed EKF-ADMM algorithm, as the correction due to the true measurements $y_{k}^{m}$ can be performed only once, instead of $n_{a}$ times, as it does not change with $t$. Moreover, there is no need for any matrix inversion when processing the fake measurements, as by processing them one by one, the matrix inversion required to compute the Kalman gain becomes a simple division, since each measurement is just a scalar value. Assuming a complexity $\mathcal{O}(n_{x})$ for evaluating the proximal operator, EKF-ADMM has complexity $\mathcal{O}(n_{x}^{3}+n_{a}n_{x}^{2})$, which is the same order of the full EKF for general state estimation. Moreover, EKF-ADMM has the same number of Jacobian matrices evaluations than the classical EKF, which is usually the most time-consuming part in case $x$ represents the weights and bias terms of a neural network model to learn. Summarizing, the proposed approach is computationally efficient and, if the Kalman filter is implemented using numerically robust factored or square-root modifications [bib16], the method is appealing for embedded applications.
@@ 4 Regret analysis
where $C_{k}$ are now given time-varying matrices for $k=0,1,\ldots,N-1$, and convex regularization terms $g$. In particular, we want to evaluate the ability of the algorithm to solve the optimization problem $\min_{x}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$ online, where $f_{k}(x)=\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x\right\rVert^{2}_{R_{k}^{-1}}$, via the following two regret functions $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(\nu_{k}))-\min_{x,\nu\in\mathcal{S}}\sum_{k=0}^{N-1}(f_{k}(x)+g(\nu))$ and $R_{c}(N)=\sum_{k=0}^{N-1}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}$, where, to simplify the notation, we have defined $x_{k}=\hat{x}_{k|k-1},P_{k}=P_{k|k-1},\forall k=0,1,\ldots,N-1$. Notice that $R_{f}(N)$ quantifies the loss we suffer by learning the model online instead of solving it in a batch way given all NN measurements, while $R_{c}(N)$ quantifies the violation of the constraint $x=\nu$. To ensure a proper behavior of EKF-ADMM, we want to prove a sublinear regret bound for both, i.e., $R_{f}(N)\leq\mathcal{O}(\sqrt{N})$ and $R_{c}(N)\leq\mathcal{O}(\sqrt{N})$ [bib2].
EKF-ADMM is a generalization of the online ADMM method proposed in [bib2], in which a sublinear regret bound is derived for the case $n_{a}=1$ and $P_{k}^{-1}=P^{-1}\succ 0,\forall k$, while, more recently, in [bib18] a sublinear regret bound has been derived for the case $n_{a}=1$ and $P_{k}^{-1}\succeq P_{k+1}^{-1},\forall k$. Here we will provide a sublinear regret in the case $n_{a}=1$ and $P_{k}^{-1}=P_{k+1}^{-1},\forall k\geq k_{n}\ll N$, which is a reasonable assumption as the EKF covariance matrix, when estimating the parameters of a model, usually has a transient and then reaches a steady-state value. By assuming $n_{a}=1$ and $P_{k}^{-1}=P_{k+1}^{-1},\forall k\geq k_{n}\ll N$, Algorithm 1 can be equivalently rewritten as in Algorithm 2.
@@
  $\begin{aligned} x_{k+1}\leftarrow&\operatorname*{arg\,min}_{x}\frac{1}{2}\left\lVert y_{k}^{m}-C_{k}x\right\rVert^{2}_{R_{k}^{-1}}+w_{k}^{T}(x-\nu_{k})+\\ &+\frac{\rho}{2}\left\lVert x-\nu_{k}\right\rVert^{2}+\frac{\eta}{2}\left\lVert x-x_{k}\right\rVert^{2}_{P_{k}^{-1}}\end{aligned}$
  $\begin{aligned} \nu_{k+1}\leftarrow&\operatorname*{arg\,min}_{\nu}g(\nu)+w_{k}^{T}(x_{k+1}-\nu)+\\ &+\frac{\rho}{2}\left\lVert x_{k+1}-\nu\right\rVert^{2}=\text{prox}_{\frac{g}{\rho}}(x_{k+1}+\frac{w_{k}}{\rho})\end{aligned}$
@@
The following Theorem 4.1 is an extension of [bib2, Theorem 4], and provides conditions for sublinear regret bounds of Algorithm 2 in the case of a linear time-varying model (12) and convex regularization function $g$.
Theorem 4.1.
@@
In general, as proved in [bib18], Theorem 4.1 holds with $M_{k_{n}}=0$ whenever $P_{k}^{-1}\succeq P_{k+1}^{-1},\forall k$. Intuitively, this means that for online model adaptation we need to limit the importance of the previous samples to promptly adapt the model to changes and therefore bound the regret function. This can be accomplished, for example, using the EKF with a proper forgetting factor [bib3].